find the common difference of the following
a. -9,-6,-3,0
b. 4,8,12,16,20
c. 6,13,20,27
d. -160,-80,-40,-20
e. 1/16,1/8,1/4.1/2
1
Expert's answer
2021-10-14T12:23:43-0400

"\text{The common difference is given by $a_{n+1}-a_n$ }\\
\text{For question a., the common difference is given by $-6-(-9)=3$}\\
\text{For question b., the common difference is given by $8-4=4$}\\
\text{For question c., the common difference is given by $13-6=7$}\\
\text{For question d., the common ratio is given by $\frac{-80}{-160}=\frac{1}{2}$}\\
\text{For question e., the common ratio is given by $\frac{\frac{1}{8}}{\frac{1}{16}}=2$}\\"